If AaBbCcDd is mated to AaBbCcDd (no gene interactions among all genes), what is the probability that the offspring will have either the phenotype aaB_CcDD or the genotype AAbbCcD_?
Answer #1

AaBbCcDd × AaBbCcDd

This is a tetrahybrid cross. Convert it to 4 monohybrid crosses.

Aa × Aa = 3/4 A_, 1/4 aa = 1/4 AA, 2/4 Aa, 1/4 aa

Bb × Bb = 3/4 B_, 1/4 bb = 1/4 BB, 2/4 Bb, 1/4 bb

Cc × Cc = 3/4 C_, 1/4 cc = 1/4 CC, 2/4 Cc, 1/4 cc

Dd × Dd = 3/4 D_, 1/4 dd = 1/4 DD, 2/4 Dd, 1/4 dd

Probability of phenotype aaB_CcDd = 1/4 × 3/4 × 2/4 × 1/4 = 6/256

Probability of genotype AAbbCcD_ = 1/4 × 1/4 × 2/4 × 3/4 = 6/256

Final answer = 6/256 + 6/256 = 12/256 = 6/128 = 3/64

Answer is 3/64

Please rate high.